OpenStudy (celinegirl):

The first four terms of a sequence are shown below: 9, 5, 1, −3 Which of the following functions best defines this sequence? A.) f(1) = 9, f(n + 1) = f(n) − 4; for n ≥ 1 B.) f(1) = 9, f(n + 1) = f(n) + 4; for n ≥ 1 C.) f(1) = 9, f(n + 1) = f(n) − 5; for n ≥ 1 D.) f(1) = 9, f(n + 1) = f(n) + 5; for n ≥ 1

OpenStudy (victoriasushchik):

f(x)_1 = 9 f(x)_2 = f(x)_1 - 4 so your sequence will be 9, 5, 1, -3, -7, -11...

OpenStudy (victoriasushchik):

so your answer will be : f(1) = 9, f( n + 1 ) = f(n) - 4, for n > 1

OpenStudy (pawanyadav):

You can check 9-5=4 5-1=4 1-(-3)=4 So every next term is 4 less than previous term.

OpenStudy (anonymous):

\[common~difference=-4\]\[f(n+1)=n-4\]yep :)

OpenStudy (victoriasushchik):

yep that's right!

OpenStudy (celinegirl):

Thank you!